# Robust Fast Adaptation from Adversarially Explicit Task Distribution Generation

### v1

@@ I.4. Distribution Adversary Implementations
Tail Risk Robustness. Due to the page limit in the main paper, we present experimental results on tail risk robustness across various confidence levels $\alpha$ for sinusoid regression and acrobot system identification. As illustrated in Figs. (14)/(15), the advantage of AR-MAML is similar to that in pendulum system identification. AR-MAML exhibits a more significant performance gain in the adversarial distribution than in the initial distribution.
@@
Impacts of Shift Distribution Constraints. Still, by varying the Lagrange multipliers $\lambda$, we include the learned task structures and the meta-testing results in Fig.s (16)/(17)/(18). For sinusoid and acrobot cases, empirical findings are similar to those in the main paper. In point robot navigation, we observe high probability density regions significantly change, and even the meta-testing results are improved in the initial task distribution with increasing $\lambda$ values.
@@
This project uses NVIDIA 1080-Ti GPUs in numeric computation. And we employ Pytorch (Paszke et al., 2019) as the deep learning toolkit in implementing experiments.

### v3

@@ I.4. Distribution Adversary Implementations
Tail Risk Robustness. Due to the page limit in the main paper, we present experimental results on tail risk robustness across various confidence levels $\alpha$ for sinusoid regression and acrobot system identification. As illustrated in Figures 14/15, the advantage of AR-MAML is similar to that in pendulum system identification. AR-MAML exhibits a more significant performance gain in the adversarial distribution than in the initial distribution.
@@
Impacts of Shift Distribution Constraints. Still, by varying the Lagrange multipliers $\lambda$, we include the learned task structures and the meta-testing results in Figures 16/17/18. For sinusoid and acrobot cases, empirical findings are similar to those in the main paper. In point robot navigation, we observe high probability density regions significantly change, and even the meta-testing results are improved in the initial task distribution with increasing $\lambda$ values.
@@
This project uses NVIDIA A100 GPUs in numeric computation. And we employ Pytorch (Paszke et al., 2019) as the deep learning toolkit in implementing experiments.